# Generating optimal Gravitational-Wave template banks with metric-preserving autoencoders

### v1

@@ 2.2 Dimensionality reduction with autoencoders
  • first, we rescale the data $c^{A}$, $A=0,\dots 9$ by the standard deviation of the first SVD component $c^{0}$ (we rescale both training and validation data, and we use half the phases in each sub-bank for training and half for validation). It is generally advisable to rescale the input data to zero mean and unit standard deviation as it helps with efficient training of the network: this is our motivation behind rescaling $c^{A}$. We call this normalization factor $\cal N$; [^2] We are aware that to avoid leakage one should define $\cal N$ using only the training data.
@@
    where it is important to emphasize the presence of the multiplicative factor $\cal N$ on the right-hand side;
  • when we train the autoencoder, we need to impose the preservation of distances between embedding and latent space, i.e. we have $\smash{L=\alpha\times L_{\rm reconstruction}+\beta\times L_{\text{distance preservation}}}$. We use the following loss functions (we denote $\smash{{\cal D}^{A}({\cal E}({d}_{\rm train}))}$ by $\smash{\tilde{d}^{A}_{\rm train}}$ here)
@@
The reconstruction error of the autoencoder is shown in Fig. 2. We see that the autoencoder performs worse than the random forest: however, we emphasize that we have not refined the training of our autoencoders after the initial pass (in view of future use in the IAS search pipeline, we trained them once for all banks and sub-banks, with the idea that they can be refined at a future time), and we trained all of them in torch.float32 precision (which can also be refined at a future time). Even so, we will see in a moment that even with this difference in reconstruction the autoencoders perform slightly better than the random forests in the effectualness tests.
Figure 4: Scatter plot of the distances in the embedding space vs. the Euclidean distances in the latent space for $1\%$ of the pairs of validation points for the three banks discussed in Section $2.$ We see that the distances are preserved extremely well in this case, at variance with the example of a unit sphere. Notice that for each of the three banks we plot only $10000$ randomly-chosen pairs of points among those having a distance $\smash{|{c}^{(ij)}_{\rm valid}|\leq 1}$ in the embedding space. This is done only to allow the reader to distinguish the three colors in the plot: including all pairs would still show the points clustering extremely close to the diagonal.
@@
In order to construct the actual templates in the autoencoder case we need to adapt the geometric placement method of Roulet:2019hzy. We first construct a uniform grid in the embedding space: since we know that the induced metric is very well approximated by ${\cal N}^{2}\delta_{\mu\nu}$, in order to achieve a spacing $\Delta c$ in the embedding space we need our grid to have spacing $\Delta e=\Delta c/{\cal N}$. The grid is centered at $\smash{{\cal E}(d^{0}=0,d^{1}=0,\dots d^{9}=0)}$: in other words, it corresponds to the mean phase across the sub-bank. Then, we proceed as in the random forest case. For each test waveform we extract its SVD components, we encode them, and we find the closest grid point in the latent space: the waveform that matches the test one is then constructed by decoding this grid point. The only remaining problem to address is how to determine which grid points represent physical waveforms and which do not. In order to do this, we train a simple normalizing flow in zuko [^4] We use an unconditional Neural Spline Flow (NSF) durkan2019neuralsplineflows with $2$ input features and $4$ layers of transforms, each transform’s internal neural network having $2$ hidden layers of $64$ neurons. Notice that the NSF spline transformations are defined over the domain [5,5][-5,5] (see here for the zuko implementation): since our features lie in this box for all banks and sub-banks we do not standardize features before training. for each bank to predict the distribution of points (the encoding of the examples we used to construct our $c^{A}$ dataset) in the latent space, and then we drop all grid points that are outside the $99.7\%$ CL contour of this distribution. The last column of Tab. 1 lists the number of templates obtained with this procedure, while Fig. 5 shows an example of the latent- and embedding-space distribution of points for Bank 0 and Bank $4$. As in the previous section, we also implement a refinement procedure by constructing a grid with a smaller grid spacing. Instead of re-running the matching procedure for all test waveforms, we consider only those waveforms whose match with a template on the un-refined grid is less than $0.95$: for these waveforms we try to find a new match on a refined grid $\Delta e\to\Delta e/2$. The effectualness test is shown in the lower panel of Fig. 3: we see that the autoencoder performs slightly better than the random forest in the case of high-mass banks, for which after refinement all test waveforms are reconstructed with an overlap larger than $0.95$ unlike the random forest case.
@@
FIG. 5a A collection of $10000$ random points within the $99.7\%$ CL contour of the distribution of latent-space points for Bank 0 and Bank $4$.
@@
FIG. 5b The first three components of the decoded $\smash{d^{A}}$ (see Eq. (4) for their definition) for the $10000$ latent-space points shown in Fig. 5a. The right panel is a zoomed-in version of the left panel.
@@
Before concluding, it is interesting to discuss the curvature of the banks. Post-facto, we can conclude that Bank 0, Bank $4$ and Bank $12$ have all negligible curvature, since for all of them we were able to find a distance-preserving coordinate chart. That is, we have seen that for all the three banks considered, the induced metric is very well approximated by ${\cal N}^{2}\delta_{\mu\nu}$. Is there a case among all our banks and sub-banks where the situation is different than this? In principle, given the high codimension of our isometric embeddings, there is no particular constraint on the size of the intrinsic curvature of our manifolds. However, we find that all our banks and sub-banks have small or negligible curvature. We discuss this further in the appendices.
3 Conclusions
In this paper we have generalized the non-linear compression of geometric-placement template banks of Wadekar:2023kym for $(2,2)$-only GW waveforms to use autoencoders instead of random forests. We have shown that, even if our autoencoders have not been refined to perform better than the random forests in predicting the phases of the GW waveforms in each bank and sub-bank – our autoencoders perform not as well as the random forests in predicting the SVD coefficients that describe the phases of waveforms in each bank and sub-bank Roulet:2019hzy – the template banks we obtain show in general comparable effectualness as those of Wadekar:2023kym, and are actually slightly more effectual at the high-mass end than those of that reference. Moreover, our template banks achieve this effectualness while having a slightly smaller number of templates than Wadekar:2023kym across the whole mass range. Future directions might include extending our method to include higher modes as in Wadekar:2023kym, and implementing these template banks in the actual IAS search pipeline.
@@ 3 Conclusions
Importantly, our geometric placement method with autoencoders relies on us being able to find a coordinate system where the metric is $\approx{\cal N}^{2}\delta_{\mu\nu}$ across the whole manifold. This might not be possible if the banks have nonzero curvature, comparable to the grid spacing $\Delta c/{\cal N}$. In this case an alternative, albeit much more complex to implement, direction one could investigate is the following:
@@
    first we train the autoencoder to reach a given level of precision in the reconstruction of the manifold;
@@
    then, since we have the possibility of computing the distance $D(e_{(i)},e_{(j)})$ between any two points on the manifold by solving the geodesic equation, we can construct for a set of random points $\{e_{(i)}\}$ the functional
@@
For this it could be beneficial, after having trained the autoencoder, to find a coordinate change $e=e(e^{\prime})$ (i.e. $\smash{c_{\rm model}\propto{\cal D}(e(e^{\prime}))}$ is the decoder now) such that the $e^{\prime}$ lie on a domain that is easier to deal with than those seen e.g. in Fig. 15. This can be achieved with the invertible transformations typically used to train normalizing flows. Even then, successfully implementing this method requires to deal with a lot of unknown variables, be it the choice of what power $\alpha$ of the inverse geodesic distance to put in Eq. (9), the fact that it is not known in advance at which distance the points will stabilize (this likely being a function of how many points we start with), and last but not least the fact that computing the geodesic distance on curved manifolds is not an easy task in general and an area of active research. [^5] Notice that, given that our metric is already in pytorch, one could try to compute $\smash{D(e_{(i)},e_{(j)})}$ by minimizing over $\Phi$ the functional $\frac{1}{2}\int_{0}^{1}{\rm d}\lambda\,g_{\mu\nu}\big(e(\lambda)_{\Phi}\big)\frac{{\rm d}e^{\mu}(\lambda)_{\Phi}}{{\rm d}\lambda}\frac{{\rm d}e^{\nu}(\lambda)_{\Phi}}{{\rm d}\lambda}\,\,$, (10) where $\smash{e^{\mu}(\lambda)_{\Phi}=e_{(j)}^{\mu}\lambda+(1-\lambda)e_{(i)}^{\mu}+\lambda(1-\lambda){\rm MLP}^{\mu}(\lambda)_{\Phi}}$. Finding the optimal weights $\Phi$ of the MLP, initializing them such that the $\rm MLP$} vanishes, will find both the geodesic curve and, from it, $\smash{D(e_{(i)},e_{(j)})}$. Knowledge of the Christoffel symbols and the formulas of Appendix A could be used to inform better the form of $\smash{e^{\mu}(\lambda)_{\Phi}}$ at the boundaries and accelerate convergence. Moreover, we notice that having the Christoffel symbols in pytorch as in Appendix A allows us to perform random geodesic walks on the manifold with a constant – albeit small – step, such that every point is at a fixed geodesic distance from the preceding one Schwarz_2023. However these points will not be distributed uniformly but according to the volume form on the manifold and, most importantly, this would not be a geometric but a stochastic placement of templates.
@@
Figure 6: Our decoder $\smash{\cal D}$ aims at reconstructing the rescaled SVD coefficients $\smash{d^{A}=c^{A}/{\cal N}}$. In order to achieve the necessary precision we first trained 55 teachers using the simple reconstruction loss $\smash{{\rm MLP}(d_{\rm train},\tilde{d}_{\rm train})}$, where $\smash{\tilde{d}^{A}_{\rm train}={\cal D}(\theta_{\rm train})}$ and $\theta_{\rm train}$ are the cosmological parameters $(\omega_{\rm c},\omega_{\rm b},n_{\rm s})$ (each batch of the training dataset contains a random selection of cosmological parameters from the regular grid used to generate the training spectra). Then, we trained a student with the same architecture as the 55 teachers using transfer learning: the loss was the sum of a reconstruction loss $\smash{{\rm MLP}(d_{\rm train},\tilde{d}^{\rm student}_{\rm train})}$ and $\smash{\smash{{\rm MLP}(\langle d^{\rm teacher}_{\rm train}\rangle,\tilde{d}^{\rm student}_{\rm train})}}$, where $\smash{\langle d^{\rm teacher}_{\rm train}\rangle}$ is the mean of the predictions of the 55 teachers. Following the initial transfer learning procedure, the student model was iteratively fine-tuned for ten subsequent iterations. All the training procedure was carried out using torch.float64 precision. The left panel shows the $99.7$th percentile of the absolute error for each of the $16$ $d^{A}$ in the validation set, for the fifth-to-last and last refinement. The right panel shows instead the $99.7$th percentile of the relative error in the validation power spectra as a function of wavenumber. By comparing the full with the dashed lines (the dashed lines showing the error with respect to the validation spectra where all the SVD components after the $16^{\rm th}$ are set to zero) shows that we are mostly dominated by the error coming from us keeping only the first $16$ basis functions.

### v2

@@ 3 Dimensionality reduction with autoencoders
  • we rescale the data $c^{A}$, $A=0,\dots 9$ by the standard deviation of the first SVD component $c^{0}$ (we rescale both training and validation data, and we use half the phases in each sub-bank for training and half for validation). It is generally advisable to rescale the input data to zero mean and unit standard deviation as it helps with efficient training of neural networks: this is our motivation behind rescaling $c^{A}$. We call this normalization factor $\cal N$.
@@
  • we train the autoencoder by imposing the preservation of distances between embedding and latent space, i.e. we have $\smash{L=\alpha\times L_{\rm reconstruction}+\beta\times L_{\text{distance preservation}}}$. We use the following loss functions (we denote $\smash{{\cal D}^{A}({\cal E}({d}_{\rm train}))}$ by $\smash{\tilde{d}^{A}_{\rm train}}$ here):
@@
    We train the autoencoders for all banks and sub-banks using $\alpha=\beta=1$ in the above loss equation. Since we are interested in preserving only local distances, we choose $\gamma=1$ (recall that the goal is to find a coordinate system where the metric is as close as possible to being flat across the whole manifold: in order to find this coordinate system – assuming that it exists – it is sufficient to impose that only the distance between neighboring points is preserved). [^1] When investigating different choices of hyperparameters for Bank 0, we tried optimizing the choice of $\gamma$ using Simulated Annealing, with an exponential cooling schedule for the temperature ($\smash{T_{k}=T_{0}\cdot\texttt{cooling\_rate}^{k}}$, with starting temperature of order of the validation loss of an initial pre-annealing run with $\gamma=1$), and a log-uniform perturbation to propose candidate $\gamma$ values ($\smash{1/\gamma_{\rm new}=1/\gamma_{\rm current}+10^{U}}$, where $U\sim\text{Uniform}({{-2},1})$). Even experimenting with different cooling rates and number of epochs for the annealing, we have not found significant improvements in the distance preservation. Hence, we fixed $\gamma=1$ throughout all our trainings for all banks and sub-banks.
@@
Figure 5: Scatter plot of the distances in the autoencoder latent space (y-axis) vs. the Euclidean distances in the SVD space (x-axis) for $1\%$ of the pairs of validation points for the three banks discussed in Section 3. We see that the distances are preserved extremely well in this case.
We construct the actual templates in the autoencoder case by adapting the geometric placement method of Roulet et al. (2019). We first construct a uniform grid in the embedding space: since we know that the induced metric is very well approximated by ${\cal N}^{2}\delta_{\mu\nu}$. In order to achieve a spacing $\Delta c$ in the embedding space, we need our grid to have spacing $\Delta e=\Delta c/{\cal N}$. The grid is centered at $\smash{{\cal E}(d^{0}=0,d^{1}=0,\dots d^{9}=0)}$, which corresponds to the mean phase across the sub-bank. We perform the effectualness test for the bank using test waveforms which are different from those used for training the model. For each test waveform, we first extract its SVD components and encode them. We then find the closest grid point in the latent space; the best-fit template to the test signal is the one obtained by decoding this grid point. We trained the autoencoders for all banks and sub-banks in torch.float32 precision (which can be refined if needed in the future).
@@
We show the reconstruction error of the autoencoder model from Eq. (3) in the lower panel of Fig. 3 and find that it outperforms the 2D SVD case discussed in the last section. The only remaining problem to address is how to determine which grid points in the autoencoder latent space represent physical waveforms and remove the unphysical ones. In order to do this, we train a simple normalizing flow in zuko [^2] We use an unconditional Neural Spline Flow (NSF) Durkan et al. (2019) with $2$ input features and $4$ layers of transforms, each transform’s internal neural network having $2$ hidden layers of $64$ neurons. Notice that the NSF spline transformations are defined over the domain [5,5][-5,5] (see here for the zuko implementation): since our features lie in this box for all banks and sub-banks we do not standardize features before training. for each bank to predict the distribution of points in the latent space (the encoding of the examples we used to construct our $c^{A}$ dataset). We then drop all grid points that are outside the $99.7\%$ CL contour of this distribution. The last column of Tab. 1 lists the number of templates obtained with this procedure, while Fig. 6 shows an example of the latent- (top panel) and SVD- (bottom panel) space distributions of points for Bank 0 and Bank $4$. We show the effectualness test of our banks in the lower panel of Fig. 4. We also show the results of a refinement procedure by constructing a new template grid in the latent space with a smaller grid spacing.
@@
FIG. 6a To visualize the latent space of the autoencoders, we show a collection of $10000$ random points within the $99.7\%$ CL contour of the distribution of latent-space points for Bank 0 and Bank $4$.
@@
FIG. 6b We project the latent-space points to the original SVD space using the decoder network to visualize the reconstructed distribution. The first three components of the decoded $\smash{d^{A}}$ (see Eq. (4) for their definition) for the $10000$ latent-space points shown in Fig. 6a. The right panel is a zoomed-in version of the left panel.
@@
Before concluding, it is interesting to discuss the curvature of the banks. Post-facto, we can conclude that Bank 0, Bank $4$ and Bank $12$ have all negligible curvature, since for all of them we were able to find a distance-preserving coordinate chart and the induced metric in each case is very well approximated by ${\cal N}^{2}\delta_{\mu\nu}$. Because our isometric embeddings have high codimension, our method does not impose specific limits on the intrinsic curvature of the manifolds we can model. However, we find that all our banks and sub-banks have small or negligible curvature. We discuss this further in Appendix E.
4 Dimensionality reduction with random forests
Let us now compare our results to the methodology of Ref. Wadekar et al. (2024b) who used a different ML algorithm for dimensionality reduction, i.e., random forest regressor (RF). We use the RF method implemented in the IAS pipeline code [^3] The IAS pipeline code is publicly available at https://github.com/JayWadekar/gwIAS-HM.. We first construct the same SVD decomposition described earlier in section 2. We use the $\rm RF$ (we use https://scikit-learn.org/stable/modules/generated/sklearn.ensemble.RandomForestRegressor.html with $50$ trees and maximum depth of $50$) to predict the function: $(c^{2},c^{3},\dots c^{9})=f(c^{0},c^{1})$. In the upper-right panel of Fig. 3, we show the phase reconstruction error using Eq. (3) in the validation dataset for this RF.
For testing the effectualness of the RF method, we again construct the templates using a uniform grid in the $(c^{0},c^{1})$ space with the grid spacing $\Delta c$ is chosen in the same way as in Roulet et al. (2019); Wadekar et al. (2024b), and listed in the third column of Tab. 1. In Fig. 4, we show the effectualness tests for the same three banks of Fig. 3. For each test waveform, we first unwrap its phase and take the dot product with the phase basis functions of Eq. (1) to find the $(c^{0}_{\rm test},c^{1}_{\rm test})$ coefficients of this test waveform. We then find the closest grid point $(c^{0}_{\rm closest},c^{1}_{\rm closest})$ in the $(c^{0},c^{1})$ space. We use this grid point to construct the best-fit template in the bank using Eq. (1), in particular, we use $(c^{0},c^{1})=(c^{0}_{\rm closest},c^{1}_{\rm closest})$ and $(c^{2},c^{3},\dots c^{9})={\rm RF}(c^{0}_{\rm closest},c^{1}_{\rm closest})$. We calculate the match of the best-fit template with the test waveform and report the result as effectualness. For reference, the top left panel of Fig. 3 shows the comparison of the case without the RF (where $(c^{2},c^{3},\dots c^{9})$ are simply put to zero), and we indeed find a large improvement due to the RF.
Similar to the previous section, we refine our effectualness calculation by matching the test waveforms to their closest grid points on a new and refined grid with spacing $\Delta c\to\Delta c/2$. Note however that, along the $(c^{0},c^{1})$ dimensions, not all the points of the rectangular grid describe physically viable waveforms. We use the method adopted in Roulet et al. (2019) to reject the grid points which do not have a physical waveform in their vicinity (see Fig. 4 of Roulet et al. (2019) for reference). The second column of Tab. 1 shows the number $\smash{N_{\rm templates}}$ of templates, i.e. how many of these grid points are “physical” for each bank and sub-bank. We find that our results are comparable to those from the autoencoder and both the machine learning algorithms outperform the linear SVD model. In Appendix A and Fig. 8, we show toy examples of cases where using autoencoders can significantly perform the RF results, however for the case of realistic GW examples, we find that both RF and autoencoder methods have comparable accuracy throughout the parameter space that we tested.
5 Discussion
@@
Our geometric placement method with autoencoders relies on us being able to find a coordinate system where the metric is equal to ${\cal N}^{2}\delta_{\mu\nu}$ across the whole manifold. This might not be physically possible if the waveform space has non-zero curvature, comparable to the grid spacing $\Delta c/{\cal N}$ (unlike the gravitational wave cases we encountered where the waveform space is nearly flat). In this case an alternative, albeit much more complex to implement, direction one could investigate is the following:
@@
    we first train the autoencoder to reach a given level of precision in the reconstruction of the manifold;
@@
    since we have the possibility of computing the distance $D(e_{(i)},e_{(j)})$ between any two points on the manifold by solving the geodesic equation, we can construct for a set of random points $\{e_{(i)}\}$ the functional
@@
It could be beneficial, after having trained the autoencoder, to find a coordinate change $e=e(e^{\prime})$ (i.e. $\smash{c_{\rm model}\propto{\cal D}(e(e^{\prime}))}$ is the decoder now) such that the $e^{\prime}$ lie on a domain that is easier to deal with than those seen e.g. in Fig. 16. This can be achieved with the invertible transformations typically used to train normalizing flows. Even then, successfully implementing this method requires us to deal with a lot of unknown variables. For example, the choice of what power $\alpha$ of the inverse geodesic distance to put in Eq. (9), and the fact that it is not known in advance at which distance the points will stabilize (this likely being a function of how many points we start with). Furthermore, the fact that computing the geodesic distance on curved manifolds is not an easy task in general and an area of active research  [^4] Notice that, given that our metric is already in pytorch, one could try to compute $\smash{D(e_{(i)},e_{(j)})}$ by minimizing over $\Phi$ the functional $\frac{1}{2}\int_{0}^{1}{\rm d}\lambda\,g_{\mu\nu}\big(e(\lambda)_{\Phi}\big)\frac{{\rm d}e^{\mu}(\lambda)_{\Phi}}{{\rm d}\lambda}\frac{{\rm d}e^{\nu}(\lambda)_{\Phi}}{{\rm d}\lambda}\,\,$, (10) where $\smash{e^{\mu}(\lambda)_{\Phi}=e_{(j)}^{\mu}\lambda+(1-\lambda)e_{(i)}^{\mu}+\lambda(1-\lambda){\rm MLP}^{\mu}(\lambda)_{\Phi}}$. Finding the optimal weights $\Phi$ of the MLP, initializing them such that the $\rm MLP$} vanishes, will find both the geodesic curve and, from it, $\smash{D(e_{(i)},e_{(j)})}$. Knowledge of the Christoffel symbols and the formulas of Appendix A could be used to inform better the form of $\smash{e^{\mu}(\lambda)_{\Phi}}$ at the boundaries and accelerate convergence. Moreover, we notice that having the Christoffel symbols in pytorch as in Appendix A allows us to perform random geodesic walks on the manifold with a constant – albeit small – step, such that every point is at a fixed geodesic distance from the preceding one Schwarz et al. (2023). However these points will not be distributed uniformly but according to the volume form on the manifold and, most importantly, this would not be a geometric but a stochastic placement of templates., we leave exploring this alternative direction to future work.
@@
Figure 7: We use our method to perform dimensionality reduction of the linear matter power spectrum in cosmology. Our decoder $\smash{\cal D}$ aims at reconstructing the rescaled SVD coefficients $\smash{d^{A}=c^{A}/{\cal N}}$. We use the knowledge distillation technique Hinton et al. (2015) to train our network. We first trained 55 teachers using the simple reconstruction loss $\smash{{\rm MLP}(d_{\rm train},\tilde{d}_{\rm train})}$, where $\smash{\tilde{d}^{A}_{\rm train}={\cal D}(\theta_{\rm train})}$ and $\theta_{\rm train}$ are the cosmological parameters $(\omega_{\rm c},\omega_{\rm b},n_{\rm s})$ (each batch of the training dataset contains a random selection of cosmological parameters from the regular grid used to generate the training spectra). Then, we trained a student with the same architecture as the 55 teachers using transfer learning: the loss was the sum of a reconstruction loss $\smash{{\rm MLP}(d_{\rm train},\tilde{d}^{\rm student}_{\rm train})}$ and $\smash{\smash{{\rm MLP}(\langle d^{\rm teacher}_{\rm train}\rangle,\tilde{d}^{\rm student}_{\rm train})}}$, where $\smash{\langle d^{\rm teacher}_{\rm train}\rangle}$ is the mean of the predictions of the 55 teachers. Following the initial transfer learning procedure, the student model was iteratively fine-tuned for ten subsequent iterations. The left panel shows the $99.7$th percentile of the absolute error for each of the $16$ $d^{A}$ in the validation set, for the fifth-to-last and last refinement. The right panel shows instead the $99.7$th percentile of the relative error in the validation power spectra as a function of wavenumber. By comparing the full with the dashed lines (the dashed lines showing the error with respect to the validation spectra where all the SVD components after the $16^{\rm th}$ are set to zero) shows that we are mostly dominated by the error coming from us keeping only the first $16$ basis functions.
5.1 Application to modeling the cosmological matter power spectrum
Let us discuss the possibility of performing dimensionality reduction using autoencoders in the space of another set of SVD coefficients, this time in the context of cosmology. More precisely, we have in mind the COBRA decomposition of the linear matter power spectrum of Ref. Bakx et al. (2025). One can generate a set of linear power spectra as a function of a number of cosmological parameters ($\omega_{\rm c}$, $\omega_{\rm b}$ and $n_{\rm s}$, for example), and construct its SVD to find a small number of basis functions with which to represent any linear power spectrum. More precisely, following Ref. Bakx et al. (2025) we generate power spectra at $500$ wavenumbers between $\smash{0.8\times 10^{-4}\,{\rm Mpc}^{-1}}$ and $\smash{4\,{\rm Mpc}^{-1}}$ on a regular ($40\times 20\times 20$) grid $(\omega_{\rm c},\omega_{\rm b},n_{\rm s})\in[0.08,0.175]\times[0.020,0.025]\times[0.8,1.2]$ using CAMB (see the last column of Tab. 1 of Bakx et al. (2025) for more details), and we extract the first $16$ coefficients ($c^{A}$) as in Ref. Bakx et al. (2025) (with the difference that besides rescaling the power spectra by the mean of the templates, we also subtract the mean of these rescaled spectra before performing the SVD): this is our training set.